OpenStudy (anonymous):

A randomly mating population of your favorite diploid animal has the dominant allele "B" and the recessive allele "b" at the frequencies 0.65 and 0.35. A sudden cataclysmic event causes all of the homozygous recessive individuals to die. You are very sad but take comfort in the fact that your favorite diploid animals are resilient creatures that simply carr on with life and get busy mating. 1) What was the percentage of homozygous recessive individuals before the cataclysm? 2) What percentage of homozygous recessive individuals born in the first generation conceived after the cataclysm?

OpenStudy (blues):

You should use the Castle-Hardy-Weinburg equation: 1 = p^2 + 2pq + q^2 where p is the frequency of the dominant allele and q is the frequency of the recessive allele. The proportion or percentage of homozygous dominant individuals equals p^2; the proportion of heterozygous individuals is 2pq and the proportion or percentage of homozygous recessive individuals is p^2. The first part of the prob just asks you to interpret that - that is, what was the proportion of homozygous recessive individuals q^2 before they were all wiped out. Is that much clear?

OpenStudy (anonymous):

Yep clear:)

OpenStudy (blues):

Cool. For the second part you will need to calculate new frequencies for p and q, since removing all the homozygous recessive individuals from the population has changed those. The CHW equation now looks like this: 1 = p^2 + 2pq. You have to solve that for p and q to obtain the new values. Then to calculate the percentage of pp individuals born in the first generation after the cataclysm, you need to recalculate 1 = p^2 + 2pq + q^2 with your new values of p and q, and use that to figure out the proportion of homozygous recessive q^2 individuals. Hope that's helpful...

OpenStudy (anonymous):

Yes it does. Thank you!